OpenStudy (anonymous):

What's next? What is the value for x in the proportion x/x - 6 = 4/7? 7x = 4(x-6) 7x = 4x - 24

OpenStudy (anonymous):

is this your equation: \[\large \frac{x}{x-6}=\frac{4}{7} \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

your last step is correct... what's wrong?

OpenStudy (anonymous):

It asks what the value for "x" is in the proportion

OpenStudy (anonymous):

not what 7x is

OpenStudy (anonymous):

on your last step, subtract 4x from both sides of the equation...

OpenStudy (anonymous):

-4x + 7x = -24

OpenStudy (anonymous):

simplify that...

OpenStudy (anonymous):

oh like terms!

OpenStudy (anonymous):

3x = -24 /3 /3

OpenStudy (anonymous):

x = -8

OpenStudy (anonymous):

good... so what's x? yes....:)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you did all the work.. good job...:)